# Leveraging Passage Embeddings for Efficient Listwise Reranking with Large Language Models

### v1

@@ 3.2 Implementation Details
As for training data, we leverage Wikipedia for alignment and MS MARCO for the learning-to-rank stage. For the latter, we use a retrieval model to obtain the top 20 candidate passages for the queries in the training set and employ a cross-encoder as the teacher model to estimate the golden ranking. More details about data construction, model selection, and implementation are listed in Appendix B and C.
During the evaluation, for each dataset, we first use a retrieval model to recall the top 100 passages for each query, and then evaluate the reranking results. For convenience, we encode the passages on the fly, allowing us to use different retrieval models to provide a more comprehensive comparison. If not otherwise specified, we use the sliding window trick to complete the whole ranking and set the window size to 20 and the step size to 10, therefore need 9 passes in total. We use one Nvidia H100 GPU to finish all evaluations.
3.3 Baselines
We select several existing methods as our basic baselines, including supervised neural reranking models monoBERT Nogueira et al. (2019) and monoT5 Nogueira et al. (2020) that are trained using a large amount of human annotation data, unsupervised LLM-based listwise approach RankGPT Sun et al. (2023), as well as several listwise ranking models that are based on smaller LLMs and trained with distillation including RankVicuna Pradeep et al. (2023a) and RankZephyr Pradeep et al. (2023b).
For a fair comparison, we train a model using a similar paradigm as RankVicuna Pradeep et al. (2023a) but use the Mistral-7B and the training data same as PE-Rank, denoted as RankMistral.
@@ 3.3 Baselines
Also, we use this model to evaluate different text compression strategies and compare them with PE-Rank. Specifically, we can use different texts to replace the original passage in the inputs, denoted as $\textbf{RankMistral}_{\ast}$, where $\ast$ can be passage ($p$), summary ($s$), or title ($t$). We provide more details on baselines in Appendix D.
4 Experiment Results
4.1 Effectiveness Analysis
@@
BM25 - 50.58 47.96 43.42
Jina-Embedding - 65.94 63.89 41.46
Supervised models trained with human annotation
monoBERT BM25 70.50 67.28 47.16
monoT5 71.83 68.89 51.36
Unspervised LLM-based listwise models
$\text{RankGPT}_{3.5}$ BM25 65.80 62.91 49.37
$\text{RankGPT}_{4}$ 75.59 70.56 53.68
LLM-based listwise models trained with distillation
RankVicuna BM25 66.82 65.49 -
RankZephyr 74.20 70.86 -
RankMistral 71.73 68.07 43.65
PE-Rank 70.48 63.54 47.96
RankVicuna Jina 69.81 70.61 -
RankZephyr 69.83 75.15 -
RankMistral 71.44 73.27 42.86
PE-Rank 70.91 69.48 44.28
Table 1: Results (NDCG@10) of reranking top-100 passages on TREC DL and BEIR. Ret means the retrieval model used in first stage.
Model $n$ TREC DL19 TREC Covid
NDCG # Proc. # Gen. Latency (s) NDCG # Proc. # Gen. Latency (s)
$\text{RankMistral}_{p}$ 20 64.65 02265.8 109.9 02.04 ($\times$.00) 70.90 08190.9 110.4 02.51 ($\times$.00)
$\text{RankMistral}_{s}$ 63.03 01490.7 106.1 01.99 ($\times$.98) 65.15 02224.2 100.2 01.92 ($\times$.76)
$\text{RankMistral}_{t}$ 48.62 00409.5 107.2 01.93 ($\times$.95) 66.71 00829.7 110.4 01.89 ($\times$.75)
PE-Rank 62.66 00326.9 020.0 00.42 ($\times$.21) 72.34 00344.3 020.0 00.44 ($\times$.18)
$\text{RankMistral}_{p}$ 100 71.96 19506.2 910.2 16.20 ($\times$.00) 77.80 71431.2 986.5 21.46 ($\times$.00)
$\text{RankMistral}_{s}$ 70.50 13485.3 881.6 15.68 ($\times$.97) 73.85 20148.6 929.6 16.94 ($\times$.79)
$\text{RankMistral}_{t}$ 45.43 03753.4 865.1 15.12 ($\times$.93) 75.40 07555.0 916.9 15.87 ($\times$.74)
PE-Rank 70.48 02942.4 180.0 03.62 ($\times$.22) 77.72 03098.9 180.0 03.65 ($\times$.17)
Table 2: Efficiency analysis for reranking top $n$ candidates retrieved by BM25 on TREC DL19 and TREC Covid. # Proc and # Gen mean the number of processed tokens in the prefill stage and generated tokens in the decode stage, respectively. For PE-Rank, we also include the time for encoding the passages on the fly. $L_{p}$ of DL19 and Covid is approximately 100 and 423, respectively. The best model in each block is in bold, and the second best is underlined.
We first evaluate the effectiveness of PE-Rank on TREC DL and BEIR benchmarks, and present the results in Table 1. From the results, we can observe that the supervised models based on BERT and T5 can achieve competitive ranking performance, while in the LLM-based baselines, using the strongest LLM, GPT-4, for listwise reranking can achieve state-of-the-art across all models on three datasets. As for distilled models, RankZephyr also shows promising ranking effectiveness, and we attribute this to using GPT-4 as the teacher model.
Comparing the proposed PE-Rank model with other baselines, we can see that: (i) without directly trained with human-annotated data, PE-Rank can approach supervised baselines’ performance. (ii) Despite compressing the entire passage into a single embedding, PE-Rank still maintains comparable effectiveness to the uncompressed distilled listwise models, even surpassing them on some datasets. For example, comparing PE-Rank with RankMistral, we can find that its ranking performance on DL19 has decreased less than 2%, while the results on BEIR are even consistently higher.
It should be emphasized that when PE-Rank remains competitive, it has a significant efficiency advantage, and we will provide a detailed analysis in the next section.
4.2 Efficiency Analysis
We conduct efficiency analysis from the perspectives of consumed tokens and latency on DL19 and Covid. We select Covid as it has a relatively long passage length, while the results on DL20 are similar to those on DL19.
Number of Consumed Tokens
We theoretically analyze the number of processed tokens in the prefill stage and generated tokens in the decode stage of different methods. Assume a single pass with $n$ passages of average length $L_{p}$ and instruction of length $L_{I}$, methods based on the text like RankGPT exhibit an input length of $O(L_{I}+nL_{p})$, which increases almost proportionally with $L_{p}$. In contrast, PE-Rank shows an input length of $O(L_{I}+n)$ which will be unchanged when $L_{p}$ increases. For RankGPT-like methods, they need to generate numbers as well as identifiers such as “[]” and may not output completely correctly, resulting in the number of generated tokens for $\Omega(mn)$. In practice $m\approx 4.5$. As for PE-Rank, by employing the DC decoding method, the number is exactly equal to $n$ since only $n$ unique special tokens will be output.
It is important to note that when employing the sliding window strategy, the above results must be multiplied by the number of sliding windows. However, PE-Rank, due to the compression of input length, can achieve completion with fewer sliding instances or even in a single pass, thereby further underscoring its efficiency advantages.
Table 2 displays the number of tokens consumed by different methods. The results show that, although simple text compression techniques partially reduce tokens to be processed, they may lead to significant performance degradation. Specifically, when using titles as compression on DL19, the performance is notably poor, possibly due to title misses or lack of valid information. Using summaries as input also results in performance loss, particularly on the Covid dataset. Besides, these text-based methods do not decrease the number of generated tokens. Note that the model may not output in the required format in practice, leading to fluctuations in the number of generated tokens.
In contrast, PE-Rank significantly reduces the number of tokens to be processed and generated, while minimizing the loss of ranking performance. Surprisingly, when ranking the top 20 passages on the Covid dataset, it even outperforms the approach without compression.
Latency
We also analyze the reranking latency using different methods in Table 2. The results indicate that heuristic text compression techniques, such as using titles or summaries, do not significantly reduce latency. Conversely, by leveraging passage embedding as a text compression representation, PE-Rank markedly accelerates the ranking process, achieving approximately a five-fold increase in speed across different candidate numbers and datasets, with only about 0.2 times the delay of the uncompressed method. Notably, when reranking the top 20 candidates, the ranking latency for a single query can be limited to 0.5 seconds, rendering it practical for real-world ranking scenarios.
Figure 4: Latency of reranking top 100 candidates at different stages during inference.
4.3 Ablation Study
Training Strategies
DL19 DL20 Covid News
(a) PE-Rank 70.48 63.54 77.72 47.40
(b) w/o Alignment 65.83 61.35 73.12 46.71
(c) w/o $\mathcal{L}_{\text{content}}$ $\mathcal{L}_{\text{KL}}$ 68.43 64.42 77.21 46.23
(d) w/o $\mathcal{L}_{\text{KL}}$ 68.43 64.03 76.33 47.42
(e) w/o $\mathcal{L}_{\text{content}}$ 66.66 60.85 75.94 47.15
Table 3: Ablation on different training strategies. We show the results of ranking top 100 candidates of BM25.
Different Embedding Models
To verify whether our proposed framework can generalize to different embedding models, we choose a different embedding model for experiments. Specifically, we select BGE-base Xiao et al. (2023), a BERT-based model that achieves the top tier position across the same parameter scale models on the MTEB benchmark Muennighoff et al. (2022). We use BGE as the embedding model and the same complete training process as Jina-Embeddings to obtain a new model. The results are shown in Table 4.
Firstly, using Jina-Embeddings and BGE as the encoder and leveraging their passage embeddings for reranking are both effective, reranking the candidates obtained from different retrieval models on different datasets can consistently bring improvement. This proves that the PE-Rank approach can be applied to different embedding models.
However, although BGE scores higher than Jina-embedding on MTEB, the performance of reranking BM25 retrieval results using BGE embeddings is consistently lower across three different datasets compared to using Jina embeddings. Due to the use of different training data and pooling methods in these two models, it is challenging to directly determine the cause of this discrepancy. Nonetheless, we have reason to believe that models excelling in general embedding benchmarks like MTEB may not necessarily perform well in this context. This issue is worth further investigation.
Model Ret. DL19 DL20 BEIR Avg.
BM25 BM25 50.58 47.96 43.80
@@ 4.3 Ablation Study
Jina-Embeddings Jina 65.94 63.89 41.46
@@
Table 4: Using different embedding models to obtain passage embeddings as context compression.
Impact of Sliding Window

### v2

@@ 4.1. Model selection
4.2. Training Data
During the alignment stage, we employ segmented Wikipedia as the training dataset. The texts in the Wikipedia dataset, authored and reviewed by humans, are of higher quality and completeness. Additionally, its comprehensive nature provides knowledge from diverse fields, rendering it reliable for training in the alignment stage. Specifically, we utilized the Wikipedia dump from Dec 2020, preprocessed by Izacard et al. (2023), totaling around 31.5 million texts. We sampled 2 million data pieces for training. The complete data format can be found in Appendix A.
In the learning-to-rank stage, we utilize the MS MARCO dataset (Bajaj et al., 2016). MS MARCO is a large-scale passage retrieval dataset that contains around 8.8 million passages and 800,000 queries, of which about 500,000 have manually annotated relevance labels. We use Jina-embeddings-v2-base-en [^2] https://huggingface.co/jinaai/jina-embeddings-v2-base-en as the retrieval model to retrieve the top 20 candidate passages for all queries in the training set, to construct the dataset. However, it only includes binary annotations (i.e., relevant or irrelevant) and cannot be directly used as training data for our training procedure. Therefore, following the approach of Zhang et al. (2023), we use an existing powerful supervised reranking model, i.e., MiniLM [^3] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2, as the annotation model to approximate the golden ranking. Following Pradeep et al. (2023a), we used a data augmentation strategy of randomly shuffling document order. To facilitate training, we excluded samples with excessively long lengths, retaining only those with input lengths less than 2048. Consequently, our dataset for this stage comprises 232,419 samples and each sample contains 20 passages and the approximated golden ranking.
4.3. Evaluation Datasets
We evaluate PE-Rank on multiple retrieval benchmarks, including TREC DL (Craswell et al., 2020) and BEIR (Thakur et al., 2021). TREC DL uses the MS MARCO dataset (Bajaj et al., 2016) as the retrieval corpus and has fine-grained relevance annotations. We use the test sets of TREC DL 2019 and TREC DL 2020, which contain 43 and 54 queries respectively. BEIR contains 18 datasets from different fields with different query requirements, aiming to evaluate the generalization ability of ranking models. Following previous work (Sun et al., 2023), we conduct evaluations on 8 datasets that contain a relatively small number of queries. We use NDCG@10 as the evaluation metric.
4.4. Baselines
We select several existing methods as our basic baselines.
Supervised Neural Rerankers
First, we select two typical supervised models, including monoBERT (Nogueira and Cho, 2019) and monoT5 (Nogueira et al., 2020). Both of these models are trained on the MS MARCO dataset using a large number of human annotation labels.
LLM-based Rerankers
Additionally, we use an unsupervised LLM-based method as the baseline: RankGPT (Sun et al., 2023), a state-of-the-art method that uses a sliding window strategy for listwise ranking based on GPT. We also add listwise reranking models that are based on smaller LLMs (such as an LLM with 7B parameters) and are distilled from RankGPT as baselines, including RankVicuna (Pradeep et al., 2023a) and RankZephyr (Pradeep et al., 2023b).
Table 1. Results (NDCG@10) of reranking top-100 passages on BEIR benchmark. Ret means the retrieval model used in the first stage. In each block, i.e., when using the same retriever, * denotes that there is no statistically significant difference between PE-Rank and the baselines ($p\geq 0.05$ level) using a two-sided t-test. The best model among all is in bold, while the best model in each block is underlined.
BM25 - 0.5947 0.3375 0.4422 0.3180 0.6789 0.3305 0.3952 0.4070 0.4380
Jina-Embeddings - 0.6894 0.3143 0.2868 0.3332 0.6553 0.2576 0.3980 0.3823 0.4146
monoBERT BM25 0.7001 0.3688 0.3175 0.4187 0.7136 0.3144 0.4462 0.4935 0.4716
monoT5 0.8071 0.3897 0.3241 0.4445 0.7657 0.3255 0.4849 0.5671 0.5136
$\text{RankGPT}_{3.5}$ 0.7667 0.3562 0.3618 0.4447 0.7043 0.3212 0.4885 0.5062 0.4937
$\text{RankGPT}_{4}$ 0.8551 0.3847 0.3857 0.4712 0.7495 0.3440 0.5289 0.5755 0.5368
RankMistral BM25 $\underline{0.7800}^{*}$ $0.3310^{*}$ $0.2746^{*}$ $0.3771^{*}$ $0.6622^{*}$ $0.3004^{*}$ 0.3710 0.3954 0.4365
PE-Rank 0.7772 0.3639 0.3306 0.4005 0.6938 0.3374 0.4970 0.4740 0.4843
RankMistral Jina $\underline{0.8019}^{*}$ $0.2974^{*}$ $0.2916^{*}$ 0.4025 $0.6385^{*}$ 0.2817 0.3580 0.3569 0.4286
PE-Rank 0.7749 0.3092 0.3000 0.3626 0.6448 0.2654 0.4478 0.4373 0.4428
However, a direct comparison with the above baseline is not intuitive because of the impact of different foundation models and training data. Furthermore, the underlying LLM and the training data are orthogonal to the ranking paradigm. Therefore, to ensure a fair comparison between the previous listwise ranking paradigm (e.g., RankGPT and RankVicuna) and PE-Rank, we retrained a listwise ranking model using the similar training process and paradigm as RankVicuna but based on the same LLM, i.e., Mistral-7B, and training data with PE-Rank, denoted as RankMistral, as a main baseline. We believe that directly comparing PE-Rank with RankMistral can provide more rich insights.
@@ 4.4. Baselines
We also use this model to evaluate different text compression strategies and compare them with PE-Rank. Specifically, we can use different texts to replace the original passage in the inputs, denoted as $\textbf{RankMistral}_{\ast}$, where $\ast$ can be passage ($p$), summary ($s$), or title ($t$). These baselines help us evaluate the effectiveness and efficiency of different compression methods under a consistent setting.
4.5. Implementation Details
Table 2. Results (NDCG@10) of reranking top-100 passages on TREC DL. Ret means the retrieval model used in the first stage. In each block, * denotes that there is no statistically significant difference between PE-Rank and the baselines ($p\geq 0.05$ level) using a two-sided t-test.
Model Ret. TREC DL19 TREC DL20
BM25 - 0.5058 0.4796
Jina-Embedding - 0.6594 0.6389
Supervised models trained with human annotation
monoBERT BM25 0.7050 0.6728
monoT5 0.7183 0.6889
Unspervised LLM-based listwise models
$\text{RankGPT}_{3.5}$ BM25 0.6580 0.6291
$\text{RankGPT}_{4}$ 0.7559 0.7056
LLM-based listwise models trained with distillation
RankVicuna BM25 $0.6682^{*}$ $0.6549^{*}$
RankZephyr 0.7420 0.7086
RankMistral $0.7173^{*}$ $0.6807^{*}$
PE-Rank 0.7048 0.6354
RankVicuna Jina $0.6981^{*}$ $0.7061^{*}$
RankZephyr $0.6983^{*}$ 0.7515
RankMistral $0.7144^{*}$ $0.7327^{*}$
PE-Rank 0.7091 0.6948
We implement all training codes based on the PyTorch framework. To optimize memory usage and accelerate training, we applied Deepspeed ZeRO stage 2 (Rasley et al., 2020) and BFloat16 mixed precision techniques. Additionally, Flash attention (Dao et al., 2022) was used to further improve training efficiency. In the alignment stage, we trained the 7B Mistral model for 1 epoch with an effective batch size of 128 and a learning rate of $1\times 10^{-4}$. In the learning-to-rank stage, we trained the model for 1 epoch with an effective batch size of 32 and a learning rate of $2\times 10^{-5}$. $\alpha$ in Eqution (8) is set to 0.2 based on prior experiments. All models were trained on 4 Nvidia H100 GPUs. The hyperparameters were determined based on empirical observations due to resource constraints.
During the evaluation, for each dataset, we first use a retrieval model to recall the top 100 passages for each query, and then evaluate the reranking results. For convenience, we encode the passages on the fly, allowing us to use different retrieval models to provide a more comprehensive comparison. If not otherwise specified, we use the sliding window trick (Sun et al., 2023) to complete the ranking and set window size to 20 and step size to 10, therefore need 9 passes in total. We use one Nvidia H100 GPU to finish all evaluations.
5. Experiment Results
5.1. Effectiveness Analysis
We first evaluate the effectiveness of PE-Rank on TREC DL and BEIR benchmarks, and present the results in Table 2 and Table 1. From the results, we can observe that the supervised models based on BERT and T5 can achieve competitive ranking performance, while in the LLM-based baselines, using the strongest LLM, GPT-4, for listwise reranking can achieve state-of-the-art across all models on three datasets. As for distilled models, RankZephyr also shows promising ranking effectiveness, and we attribute this to using GPT-4 as the teacher model.
Table 3. Efficiency analysis for reranking top $n$ candidates retrieved by BM25 on TREC DL19 and Covid. # Proc and # Gen mean the number of processed tokens in the prefill stage and generated tokens in the decode stage, respectively. For PE-Rank, we also include the time for encoding the passages on the fly. $L_{p}$ of DL19 and Covid is approximately 100 and 423. In each block, * denotes no statistically significant difference between the compression settings and $\text{RankMistral}_{p}$ ($p\geq 0.05$ level) using a two-sided t-test. The subscript of RankMistral means the form of inputs, including original passage ($p$), summary ($s$), or title ($t$).
Model $n$ TREC DL19 Covid
NDCG@10 # Proc. # Gen. Latency (s) NDCG@10 # Proc. # Gen. Latency (s)
$\text{RankMistral}_{p}$ 20 0.6465 02265.8 109.9 02.04 ($\times$.00) 0.7090 08190.9 110.4 02.51 ($\times$.00)
$\text{RankMistral}_{s}$ $\underline{0.6303}^{*}$ 01490.7 106.1 01.99 ($\times$.98) 0.6515 02224.2 100.2 01.92 ($\times$.76)
$\text{RankMistral}_{t}$ 0.4862 00409.5 107.2 01.93 ($\times$.95) $0.6671^{*}$ 00829.7 110.4 01.89 ($\times$.75)
PE-Rank $0.6266^{*}$ 00326.9 020.0 00.42 ($\times$.21) $\textbf{0.7234}^{*}$ 00344.3 020.0 00.44 ($\times$.18)
$\text{RankMistral}_{p}$ 100 0.7196 19506.2 910.2 16.20 ($\times$.00) 0.7780 71431.2 986.5 21.46 ($\times$.00)
$\text{RankMistral}_{s}$ $\underline{0.7050}^{*}$ 13485.3 881.6 15.68 ($\times$.97) $0.7385^{*}$ 20148.6 929.6 16.94 ($\times$.79)
$\text{RankMistral}_{t}$ 0.4543 03753.4 865.1 15.12 ($\times$.93) $0.7540^{*}$ 07555.0 916.9 15.87 ($\times$.74)
PE-Rank $0.7048^{*}$ 02942.4 180.0 03.62 ($\times$.22) $\underline{0.7772}^{*}$ 03098.9 180.0 03.65 ($\times$.17)
Comparing the proposed PE-Rank model with other baselines, we can see that: (i) PE-Rank can approach supervised baselines’ performance. (ii) Despite compressing the entire passage into a single embedding, PE-Rank maintains comparable results to the uncompressed distilled listwise models, especially RankMistral. Specifically, we can find that the ranking performance of PE-Rank on both DL19 and DL20 has no statistically significant difference compared with RankMistral. On BEIR, there is also no significant difference on most datasets, even on some datasets PE-Rank surpassing RankMistral. This observation indicates that under the same settings, i.e., the same LLM and training data, PE-Rank can achieve comparable effectiveness to the previous listwise ranking paradigm.
While PE-Rank remains competitive, it has significant efficiency advantages. We provide a detailed analysis in the next section.
5.2. Efficiency Analysis
We conduct efficiency analysis from the perspectives of consumed tokens and latency. Here we conduct experiments on the TREC DL19 dataset and one of the datasets of BEIR, i.e., Covid. TREC DL19 and DL20 have the same corpus and similar distribution, therefore we only show the results on one. Instead, we select the Covid dataset as an alternative since it has longer documents.
Figure 4. Latency of reranking top 100 candidates at different stages during inference. $p$ (passage), $s$ (summary), and $t$ (title) denote the different forms of inputs of RankMistral.
Number of Consumed Tokens. We theoretically analyze the number of processed tokens in the prefill stage and generated tokens in the decode stage of different methods. Assume a single pass with $n$ passages of average length $L_{p}$ and instruction of length $L_{I}$, methods based on the text like RankGPT exhibit an input length of $O(L_{I}+nL_{p})$, which increases almost proportionally with $L_{p}$. In contrast, PE-Rank shows an input length of $O(L_{I}+n)$ which will be unchanged when $L_{p}$ increases. For RankGPT-like methods, they need to generate numbers as well as identifiers such as “[]” and may not output completely correctly, resulting in the number of generated tokens for $\Omega(mn)$. In practice $m\approx 4.5$. As for PE-Rank, by employing the DC decoding method, the number is exactly equal to $n$ since only $n$ unique special tokens will be output.
It is important to note that when employing the sliding window strategy, the above results must be multiplied by the times of sliding. However, PE-Rank, due to the compression of input length, can achieve completion with fewer times or even in a single pass, thereby further underscoring its efficiency advantages.
Table 3 displays the number of tokens consumed by different methods. The results show that, although simple text compression techniques partially reduce tokens to be processed, they may lead to performance degradation. Specifically, when using titles as compression on DL19, i.e., $\text{RankMistral}_{t}$, the performance is even lower than BM25, possibly due to title misses or lack of valid information. Using summaries as input also results in performance loss, particularly on the Covid dataset. Besides, these text-based methods do not decrease the number of generated tokens. Note that the model may not output in the required format in practice, leading to fluctuations in the number of generated tokens. In contrast, PE-Rank significantly reduces the number of tokens to be processed and generated, while there is no statistically significant difference compared with $\text{RankMistral}_{p}$ on all datasets and reranking settings.
Latency. We also analyze the reranking latency using different methods in Table 3. The results indicate that heuristic text compression techniques, such as using titles or summaries, do not significantly reduce latency. Conversely, by leveraging passage embedding as a text compression representation, PE-Rank markedly accelerates the ranking process, achieving approximately a five-fold increase in speed across different candidate numbers and datasets, with only about 0.2 times the delay of the uncompressed method. Notably, when reranking the top 20 candidates, the ranking latency for a single query can be reduced to 0.5 seconds, which for the first time makes the LLM-based listwise ranking method practical for being employed in an online search system.
5.3. Ablation Study
5.3.1. Training Strategies
Table 4. Ablation on different training strategies. We show the results of ranking top 100 candidates of BM25.
DL19 DL20 Covid News
(a) PE-Rank 0.7048 0.6354 0.7772 0.4740
(b) w/o Alignment 0.6583 0.6135 0.7312 0.4671
(c) w/o $\mathcal{L}_{\text{content}}$ $\mathcal{L}_{\text{KL}}$ 0.6843 0.6442 0.7721 0.4623
(d) w/o $\mathcal{L}_{\text{KL}}$ 0.6843 0.6403 0.7633 0.4742
(e) w/o $\mathcal{L}_{\text{content}}$ 0.6666 0.6085 0.7594 0.4715
5.3.2. Different Embedding Models
To verify whether our proposed framework can generalize to different embedding models, we choose a different embedding model for experiments. Specifically, we select BGE-base (Xiao et al., 2023), a BERT-based model that achieves the top tier position across the same parameter scale models on the MTEB benchmark (Muennighoff et al., 2022). We use BGE as the embedding model and the same complete training process as Jina-Embeddings to obtain a new model. The results are shown in Table 5.
Firstly, using Jina-Embeddings and BGE as the encoder and leveraging their passage embeddings for reranking are both effective, reranking the candidates obtained from different retrieval models on different datasets can consistently bring improvement. This demonstrates that the PE-Rank approach can be applied to different embedding models.
However, although BGE scores higher than Jina-embedding on MTEB, the performance of reranking BM25 retrieval results using BGE embeddings is consistently lower across three different datasets compared to using Jina embeddings. Due to the use of different training data and pooling methods in these two models, it is challenging to directly determine the cause of this discrepancy. Nonetheless, we have reason to believe that models excelling in general embedding benchmarks may not necessarily perform well in this context. This issue is worth further investigation.
Table 5. Using different embedding models to obtain passage embeddings as context compression.
@@ 5.3.2. Different Embedding Models
BM25 BM25 0.5058 0.4796 0.4380
@@
Jina-Embeddings Jina 0.6594 0.6389 0.4146
@@
5.3.3. Impact of Sliding Window